OpenStudy (anonymous):

In a triangle, the measure of the first angle is twice the measure of the second angle. The measure of the 3rd angle is 68 degrees more than the measure of the 2nd angle. Use the fact that the sum of the measures of the 3 angles of a triangle is 180 degrees to find the measure of each angle. The measure of the smallest angle is??

OpenStudy (ajprincess):

Let us assume the scond angle to be x. Then first angle=2x Third angle=x+68 x+2x+x+68=180 Can u find x nw?

OpenStudy (anonymous):

ok so im going to combine like terms right?

OpenStudy (ajprincess):

ya. that's the first step.

OpenStudy (anonymous):

ok so i get 4x+68=180. so now i subtract 68 from both sides?

OpenStudy (ajprincess):

ya

OpenStudy (anonymous):

then i get 4x=112 so i divide by 4 on both sides and get 28? After that how do i find the measure of the smallest angle?

OpenStudy (ajprincess):

ya

OpenStudy (anonymous):

Ok so how do i find the measure of the smallest angle?

OpenStudy (ajprincess):

The value of x is the measure of small angle.

OpenStudy (anonymous):

so 28 is the measure of the smallest angle?

OpenStudy (ajprincess):

ya.:)

OpenStudy (anonymous):

what would the measure of the 2nd largest angle be?

OpenStudy (ajprincess):

2x

OpenStudy (anonymous):

that's not right it said.

OpenStudy (anonymous):

It says what would the measure of the 2nd largest angle be?

OpenStudy (ajprincess):

2x=?

OpenStudy (anonymous):

what do i multiply there where that x is ?

OpenStudy (anonymous):

would i do 2(28)?

OpenStudy (ajprincess):

yes

OpenStudy (anonymous):

and how would i find the measure of the largest angle?

OpenStudy (ajprincess):

x+68=?

OpenStudy (anonymous):

so 28+68?

OpenStudy (anonymous):

so would the largest angle be 96?

OpenStudy (ajprincess):

yes

OpenStudy (anonymous):

If i post another one kind of similar will u help me on it? you are great

OpenStudy (ajprincess):

ya sure. post it.